OpenStudy (anonymous):

Stats Question: A multiple-choice quiz has 15 questions, each with 4 possible answers of which only 1 is the correct answer. What is the probability that sheer guesswork yields from 5 to 10 correct answers? Please help me to solve this question with explanation. Shami Shami_Sweet@yahoo.com

Directrix (directrix):

@Shami Have you studied binomial probabilty?

OpenStudy (anonymous):

Yes i did the same . I have solved many question but i am unable to understand this

OpenStudy (anonymous):

Its like : nCx PxQ n-x

Directrix (directrix):

Then, when I write that the probability of getting *exactly 5 questions correct* by guessing is this: P(5) = C(15,5) * (.25) ^5 * (.75) ^10 then do you get where those numbers come from? @Shami

OpenStudy (anonymous):

I think answers is 0.165145981

Directrix (directrix):

C (15, 5) means a combination of 15 questions taken 5 at the time and gives the number of ways groups of 5 questions can be selected from the 15 questions.

Directrix (directrix):

>>answers is 0.165145981 Answer to what? The entire question?

OpenStudy (anonymous):

great dear.. i got

OpenStudy (anonymous):

Not its answer is for P(X=5)

Directrix (directrix):

I am not interested in the answer at this point. I am interested in whether or not you know enough yet to follow how to do this problem after it we get it set up. Then, you can do the calculations. So, if you just want the answer, then I can't help you because I don't know the answer but I do know how to find it.

OpenStudy (anonymous):

Yes i dont need answers, i just to find out calculation of whole question

Directrix (directrix):

To do the problem, the individual probabilities for P(5), P(6), P(7), P(8), P(9), and P(10) will be calculated and then added to answer this question: What is the probability that sheer guesswork yields from 5 to 10 correct answers?

OpenStudy (anonymous):

okay. i try to do it

Directrix (directrix):

P(5) = C(15,5) * (.25) ^5 * (.75) ^10 + P(6) = C(15,6) * (.25) ^6 * (.75) ^ 9 + P(7) = C(15,7) * (.25)^7 * (.75) ^ 8 + P(8) = C(15,8) * (.25)^8 * (.75) ^ 7 + P(9) = C(15,9) * (.25)^9 * (.75) ^ 6 + P(10)=C(15,10) *(.25)^10 (.75) ^ 5 = @Shami Crank out those numbers and add. I'll do the same and we can compare, okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

my answer is 0.3133846

Directrix (directrix):

.31336 approx -- I truncated some decimals. I'd say that we have the same answer.

Directrix (directrix):

Check out the online binomial probability calculator at: http://easycalculation.com/statistics/binomial-distribution.php

OpenStudy (anonymous):

Thank you so much dear. M obliged really.

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

Find the surface area of a square pyramid with a base that measures 10 m on each side and has a slant height of 20 m.

OpenStudy (anonymous):

2*10*15 + 10^2 this is what I did. and got 400.

Directrix (directrix):

2*10*20 + 10^2 = ? The slant height is 20, not 15.

Directrix (directrix):

@dopegyal_ What did you get?

OpenStudy (anonymous):

Oh wow... I need sleep . lol . 500 ?(:

Directrix (directrix):

Right. 500 is what I got.